if the length of a rectangle is 3 inches longer than the width and the perimeter of the rectangle is 78. what is the equation involving only the length? also solve the equation finding the length and the width.
Answer by ptaylor(2198)

Let l-Length of rectangle
Length(l)=Width(w)+3 subtract 3 from both sides
l-3=w+3-3 or
w=l-3
Now we know that the perimeter =2l+2w or
78=2l+2(l-3) get rid of parens
78=2l+2l-6 add 6 to both sides
78+6=2l+2l-6+6 collect like terms
4l=84 divide both sides by 4
l=21-----------------------------------length of rectangle
w=l-3=21-3=18----------------------------width of rectangle
CK
2*21+2*18=78
42+36=78
78=78
Length is 3" longer than width
21 is 3 more than 18

Hope this helps----ptaylor